# TimeBench: A Comprehensive Evaluation of Temporal Reasoning Abilities in Large Language Models

### v1

@@ A.4 Prompts
TimeXNLI Date Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE
s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM E-F1 H-EM H-F1 S-EM S-F1 O-EM O-F1 C-EM C-F1 L2-EM L2-F1 L3-EM L3-F1 Acc Sym. Comm. Event Overall
Human 98 96 92 100 64 80.8 75.8 87.1 97.8 - 100 89 93.3 87 91.1 82 85.6 84 87.3 76 79.9 96 97.1 94 95.3 82.5 96.5 84.3 88.0 88.5
GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 60.6 40.4 46.5 44.4 57.0 49.0 57.0 22.0 23.1 91.0 95.3 94.0 95.0 64.8 75.8 65.4 59.3 63.5
  + CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 61.3 33.0 41.2 43.4 54.6 53.0 59.6 20.0 22.6 93.0 97.0 93.0 94.5 58.0 77.0 66.2 58.3 63.2
  + FS 85.3 73.3 53.3 100 51.0 64.8 77.5 88.3 85.0 94.6 88.6 59.2 73.7 40.0 51.0 59.6 72.4 48.0 54.8 25.3 28.7 86.0 92.4 94.8 95.9 62.8 78.0 78.6 63.0 69.5
  + FS/CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 66.9 44.4 52.8 48.5 65.3 44.0 52.6 22.0 25.9 91.0 96.9 93.0 94.6 66.4 85.0 68.3 60.8 66.5
GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.5 39.2 69.1 62.3 60.5 70.8 29.5 35.4 36.5 40.9 37.5 43.9 21.0 22.9 73.6 81.2 61.8 73.8 57.4 60.3 49.0 49.8 51.2
  + CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 64.4 29.0 35.1 35.8 39.7 38.5 42.9 24.0 26.3 32.0 57.6 54.2 68.0 52.0 50.8 35.1 43.5 42.6
  + FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 66.1 37.9 48.4 37.8 43.2 43.5 51.6 16.0 17.9 77.7 84.7 70.0 78.0 55.0 53.9 61.2 52.1 54.8
  + FS/CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 68.0 37.5 47.0 38.1 42.5 37.5 41.7 33.0 37.8 86.2 89.9 68.0 76.6 50.2 61.1 40.2 54.0 51.8
Baichuan2 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 55.4 29.0 34.6 31.0 48.8 34.0 44.3 30.0 39.5 40.0 57.4 45.0 61.4 49.0 37.9 37.5 42.9 40.7
  + CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 41.5 36.0 40.9 39.0 52.0 27.0 38.5 29.0 43.2 46.0 62.8 46.0 64.3 55.0 34.5 33.5 43.8 39.8
  + FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 60.7 35.0 45.7 37.0 51.9 31.0 41.5 19.0 31.8 73.0 81.1 48.0 59.4 48.0 46.1 46.4 47.3 46.9
  + FS/CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 57.8 27.0 36.7 38.0 49.8 34.0 40.7 33.0 43.0 72.8 80.4 43.0 60.2 44.0 48.5 36.8 46.9 44.7
LLaMA2 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 61.9 21.0 30.5 28.0 46.1 23.0 36.1 18.0 26.9 43.0 53.1 55.0 69.4 49.0 33.9 27.8 40.5 36.1
  + CoT 36.0 50.0 38.0 6.0 7.3 39.1 14.0 51.7 10.0 36.9 - 45.0 58.7 30.0 38.9 20.0 40.9 18.0 32.5 21.0 33.6 43.0 58.0 56.0 68.4 47.0 32.5 26.5 40.7 36.0
  + FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.5 15.0 62.3 40.2 24.0 34.2 17.0 18.4 11.0 25.9 5.0 14.6 22.0 33.3 54.0 68.1 50.0 64.8 47.0 45.1 35.4 32.6 35.3
  + FS/CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 46.0 21.0 25.4 34.0 46.7 23.0 36.5 7.0 16.5 72.0 80.8 54.0 66.2 50.0 43.8 28.4 40.9 38.4
Vicuna-1.5 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 60.4 29.0 37.0 38.0 46.8 22.0 37.4 17.0 23.2 14.0 42.1 13.0 43.6 46.0 34.0 31.5 34.2 33.4
  + CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 56.9 31.0 36.4 16.0 38.2 25.0 37.7 13.0 20.4 31.0 49.0 29.0 49.1 51.0 33.3 22.2 35.2 31.8
  + FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 58.3 23.0 25.9 38.0 42.6 26.0 41.4 18.0 20.1 51.0 61.8 28.0 42.6 56.0 43.4 31.1 38.5 37.3
  + FS/CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 59.5 27.0 30.7 39.0 48.1 31.0 35.9 26.0 31.2 71.0 77.5 53.0 65.5 52.0 43.9 26.9 46.3 41.3
FLAN-T5 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 61.7 24.0 26.8 31.0 33.6 48.0 52.2 20.0 21.8 84.0 87.9 78.0 83.9 64.0 39.8 35.0 51.5 45.3
  + CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 64.4 23.9 28.2 39.0 41.6 46.0 50.2 28.0 30.6 73.0 79.5 57.0 68.9 55.0 41.8 31.3 49.5 43.9
  + FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 64.3 23.0 25.0 31.0 33.6 47.0 50.6 21.0 22.4 82.0 87.0 78.0 84.5 65.0 41.1 34.2 51.3 45.1
  + FS/CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 59.4 19.0 21.2 34.0 36.6 45.0 49.2 20.0 21.7 89.0 93.8 72.0 79.7 66.0 42.9 31.1 50.7 44.7
Mistral 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 28.2 5.0 21.4 4.0 24.3 7.0 22.3 4.0 21.7 2.0 39.6 1.0 31.6 51.0 41.6 31.8 18.0 25.3
  + CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 31.3 11.0 22.4 8.0 21.1 14.0 24.9 12.0 25.6 5.0 34.0 4.0 31.2 61.0 36.4 23.5 21.2 24.2
  + FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 43.5 12.0 23.9 4.0 21.3 7.0 21.2 7.0 23.0 23.0 48.9 23.0 44.9 57.0 40.3 41.9 25.6 32.2
  + FS/CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.1 8.0 34.4 - 13.0 33.7 11.0 25.3 9.0 26.4 9.0 24.3 9.0 20.3 68.0 78.6 42.0 57.4 50.0 37.9 27.0 31.8 31.6
ChatGLM3 38.0 50.0 34.0 2.0 3.0 34.1 7.0 43.6 14.0 56.7 38.9 20.0 41.2 14.0 31.7 25.0 33.8 17.0 26.0 24.0 32.2 42.0 57.0 30.0 54.0 50.0 31.0 28.2 33.2 31.5
  + CoT 27.0 49.0 37.0 0.0 1.0 24.8 3.0 37.1 10.0 44.8 - 24.0 41.7 10.0 25.4 27.0 34.6 22.0 28.1 35.0 41.2 28.0 44.5 27.0 52.0 48.0 28.3 20.1 32.6 28.9
  + FS 37.0 52.0 30.0 0.0 10.0 53.0 9.0 52.9 19.0 54.7 50.1 11.0 19.6 13.0 17.0 25.0 35.0 24.0 30.4 20.0 25.6 27.0 46.5 33.0 53.3 54.0 29.8 35.5 29.0 30.9
  + FS/CoT 32.0 66.0 31.0 0.0 4.0 34.8 8.0 43.6 11.0 44.0 - 25.0 43.3 19.0 27.5 23.0 30.2 19.0 25.7 36.0 43.0 50.0 62.8 34.0 56.0 48.0 32.3 24.2 36.2 32.7
Baichuan2-7B 35.0 50.0 37.0 4.5 4.0 47.9 10.5 55.3 15.0 54.3 42.0 26.0 41.5 20.0 34.7 20.0 35.1 19.0 31.2 6.0 20.4 22.0 43.4 29.0 47.7 55.0 31.6 32.7 30.1 31.0
  + CoT 38.0 43.0 32.0 1.0 5.3 37.9 13.0 58.0 15.0 44.2 - 41.0 53.5 28.0 38.8 29.0 39.9 23.0 33.2 18.0 29.3 21.0 41.2 29.0 47.2 54.0 28.5 28.9 35.1 32.5
  + FS 40.0 50.0 36.0 20.0 28.7 59.4 26.5 66.9 17.0 53.0 49.8 45.0 60.7 30.0 42.1 27.0 37.8 23.0 35.7 10.0 20.4 40.0 57.4 37.0 53.0 51.0 36.5 43.0 38.0 39.1
  + FS/CoT 41.0 50.0 36.0 23.5 13.0 45.7 17.5 58.1 7.0 39.2 - 36.0 51.2 29.0 43.0 42.0 52.5 25.0 39.3 20.0 31.0 57.0 70.1 39.0 60.2 49.0 37.6 30.1 43.0 39.0
LLaMA2-7B 39.0 53.0 30.0 13.0 2.7 39.3 4.0 41.0 1.0 6.3 24.5 37.0 49.0 14.0 29.0 7.0 26.8 8.0 21.1 9.0 16.0 48.0 63.9 32.0 47.9 49.0 33.8 17.0 30.5 27.4
  + CoT 44.0 50.0 33.0 5.0 2.7 35.0 4.5 40.0 1.0 1.7 - 27.0 49.9 17.0 31.6 11.0 31.4 10.0 24.5 7.0 17.8 44.0 56.9 32.0 48.1 46.0 33.0 14.1 30.3 26.8
  + FS 44.0 60.0 34.0 11.0 4.0 62.8 8.0 64.7 8.0 40.0 30.5 36.0 50.8 20.0 29.4 5.0 22.3 6.0 18.0 6.0 17.9 12.0 36.3 23.0 44.3 53.0 37.3 31.1 25.3 28.7
  + FS/CoT 38.0 51.0 36.0 14.5 11.0 42.8 25.0 65.6 13.0 53.4 - 36.0 53.5 21.0 34.1 1.0 13.6 3.0 11.2 5.0 14.0 22.0 46.7 21.0 42.3 51.0 34.9 35.1 25.0 29.0
Vicuna-1.5-7B 37.0 58.0 43.0 5.0 1.3 40.4 9.5 52.5 6.0 32.0 47.8 35.0 47.1 11.0 18.5 20.0 35.7 15.0 25.7 12.0 17.3 14.0 33.0 14.0 46.8 54.0 35.8 27.1 26.6 28.1
  + CoT 36.0 50.0 36.0 1.5 1.3 39.4 8.5 49.2 9.0 36.2 - 30.0 40.9 14.0 24.6 16.0 26.2 14.0 28.5 12.0 25.0 9.0 27.7 7.0 40.3 54.0 30.9 23.9 24.6 25.5
  + FS 43.0 57.0 37.0 8.5 3.3 44.6 5.5 42.1 7.0 36.8 67.0 24.0 31.9 12.0 14.9 16.0 21.8 20.0 27.5 17.0 22.2 13.0 34.3 6.0 32.2 54.0 36.4 29.5 23.1 26.9
  + FS/CoT 35.0 54.0 35.0 8.0 2.7 37.2 10.0 47.5 10.0 41.3 - 31.0 39.9 13.0 16.6 15.0 26.7 15.0 23.8 16.0 23.1 55.0 66.3 32.0 48.1 43.0 33.0 24.8 31.0 29.8
${}^{\dagger}$ 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 59.4 40.0 54.2 42.0 52.7 31.0 38.0 13.0 21.4 68.0 77.3 50.0 63.5 54.0 40.4 44.6 47.6 45.7
  + CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 58.2 39.0 49.6 39.0 49.8 34.0 40.1 37.0 45.6 73.0 81.3 46.0 65.6 60.0 46.8 39.8 50.9 47.6
${}^{\dagger}$ 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 59.3 34.0 48.6 41.0 49.6 38.0 43.4 34.0 37.5 68.0 78.7 49.0 62.7 58.0 40.9 43.7 50.0 46.9
  + CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 59.4 37.0 49.1 49.0 58.4 34.0 43.8 38.0 44.1 70.0 78.0 55.0 68.2 58.0 47.5 38.5 52.5 48.4
${}^{\dagger}$ 27.0 66.0 41.0 32.5 28.0 59.8 34.5 69.4 5.0 34.3 59.8 40.0 53.8 35.0 50.1 41.0 49.6 33.0 38.5 18.0 22.9 49.0 65.9 34.0 51.0 55.0 41.6 41.5 42.4 42.1
  + CoT 30.0 56.0 34.0 34.0 23.3 57.0 33.0 69.5 12.0 44.5 - 41.0 51.2 31.0 40.7 38.0 46.4 26.0 32.6 41.7 46.3 46.0 61.5 43.8 64.1 53.0 38.5 39.9 44.2 42.3
${}^{\dagger}$ 26.0 50.0 30.0 20.0 19.3 54.5 20.0 59.6 15.0 45.2 62.4 44.0 54.4 30.0 45.3 42.0 49.8 34.0 41.9 30.0 35.8 50.0 64.0 36.0 53.3 49.0 31.5 39.4 44.0 40.8
  + CoT 37.0 52.0 36.0 25.5 21.3 56.9 26.5 67.0 16.0 41.9 - 32.0 45.6 27.0 36.1 41.0 50.9 30.0 38.0 51.0 57.3 45.0 59.7 37.0 57.7 50.0 37.6 38.3 43.9 41.5
${}^{\dagger}$ 48.0 53.0 38.0 41.0 34.0 61.8 42.5 76.2 35.0 61.8 58.3 43.0 55.9 30.0 45.3 37.0 49.4 38.0 47.8 37.0 45.5 68.0 76.7 64.0 74.8 53.0 45.0 52.8 51.0 50.6
  + CoT 57.0 63.0 35.0 54.0 30.0 61.8 42.0 45.7 29.0 57.3 - 51.0 60.4 30.0 46.2 48.0 57.2 37.0 47.9 24.0 33.2 60.0 65.9 58.0 67.9 57.0 52.3 44.3 49.6 48.7
${}^{\dagger}$ 48.0 70.0 32.0 35.0 3.3 51.8 13.5 62.6 11.0 55.0 61.6 50.0 57.2 24.0 26.3 30.0 35.4 38.0 41.5 22.0 22.5 67.0 76.4 35 55.9 58.0 46.3 37.0 42.6 41.6
  + CoT 47.0 68.0 32.0 46.0 8.7 53.9 15.5 64.3 13.0 56.5 45.0 52.5 23.0 24.5 30.0 35.0 37.0 40.2 22.0 22.5 72.0 79.4 42 60.3 54.0 48.3 35.3 42.6 41.8
Method BLEU-4 METEOR ROUGE-L CIDEr MATCH Overall Norm
[…]

### v2

@@ B.3 Prompts
TimeXNLI Date Arith DurationQA McTACO TimeDial SitGen TimeQA MenatQA TempReason TRACIE Sym. Comm. Event Avg.
s1 s2 s3 Acc EM F1 EM F1 EM F1 Norm E-EM E-F1 H-EM H-F1 S-EM S-F1 O-EM O-F1 C-EM C-F1 L2-EM L2-F1 L3-EM L3-F1 Acc
Human 98.0 96.0 92.0 100.0 64.0 80.8 75.8 87.1 97.8 97.8 100.0 89.0 93.3 87.0 91.1 82.0 85.6 84.0 87.3 76.0 79.9 96.0 97.1 94.0 95.3 82.5 96.5 91.4 89.0 91.5
GPT-4 78.6 76.0 50.7 98.0 35.0 59.2 61.2 80.0 72.0 91.1 59.3 48.9 60.6 40.4 46.5 44.4 57.0 49.0 57.0 22.0 23.1 91.0 95.3 94.0 95.0 64.8 75.8 72.4 62.4 68.3
+ CoT 80.0 76.0 60.0 92.0 35.0 58.1 67.0 82.6 65.0 89.3 - 50.0 61.3 33.0 41.2 43.4 54.6 53.0 59.6 20.0 22.6 93.0 97.0 93.0 94.5 58.0 77.0 76.7 61.1 68.5
+ FS 85.3 73.3 53.3 100.0 51.0 64.8 77.6 88.3 85.0 94.6 88.6 59.2 73.7 40.0 51.0 59.6 72.4 48.0 54.8 25.3 28.7 86.0 92.4 94.8 95.9 62.8 78.0 84.1 66.5 73.7
+ FS CoT 92.0 84.0 64.0 100.0 42.0 55.1 68.0 72.3 79.0 93.4 - 48.0 66.9 44.4 52.8 48.5 65.3 44.0 52.6 22.0 25.9 91.0 96.9 93.0 94.6 66.4 85.0 73.6 65.2 72.1
GPT-3.5 45.4 67.6 31.2 97.0 19.2 50.5 34.1 68.6 39.2 69.1 62.3 60.5 70.8 29.5 35.4 36.5 40.9 37.5 43.9 21.0 22.9 73.6 81.2 61.8 73.8 57.4 60.3 62.6 53.3 57.4
+ CoT 33.6 64.8 33.6 71.0 12.4 23.2 28.1 45.1 34.6 67.0 - 52.5 64.4 29.0 35.1 35.8 39.7 38.5 42.9 24.0 26.3 32.0 57.6 54.2 68.1 52.0 50.8 45.1 48.3 48.3
+ FS 52.0 68.4 31.6 63.6 42.8 67.7 43.5 71.2 47.8 76.4 79.1 53.8 66.1 37.9 48.4 37.8 43.2 43.5 51.6 16.0 17.9 77.7 84.7 70.0 78.0 55.0 53.9 73.6 55.6 59.7
+ FS CoT 51.6 71.8 36.6 84.4 20.8 41.2 21.4 38.1 48.3 71.1 - 56.5 68.0 37.5 47.0 38.1 42.5 37.5 41.7 33.0 37.8 86.2 89.9 68.0 76.6 50.2 61.1 50.1 56.7 56.6
LLaMA270b 44.0 47.0 32.0 78.5 12.7 59.2 23.0 68.9 10.0 57.0 25.0 28.0 40.8 31.0 40.6 8.0 18.9 11.0 16.6 9.0 12.0 50.0 63.5 39.0 54.5 48.0 50.4 52.5 36.8 44.1
+ CoT 30.0 66.0 28.0 53.5 8.0 57.3 21.0 67.1 9.0 58.6 17.0 31.4 13.0 19.5 5.0 12.2 8.0 12.7 18.0 20.8 12.0 37.5 20.0 40.5 51.0 44.4 61.0 28.2 39.1
+ FS 49.0 42.0 38.0 62.0 1.3 61.2 13.0 66.5 6.0 56.6 62.8 41.0 51.1 16.0 20.0 8.0 16.4 17.0 19.9 18.0 18.7 34.0 52.2 31.0 41.1 51.0 47.8 61.8 33.8 44.3
+ FS CoT 54.0 63.0 40.0 69.5 8.0 55.2 21.5 62.1 6.0 56.4 36.6 50.9 34.0 42.4 28.0 38.6 19.0 29.3 18.0 21.9 77.0 83.1 65.0 74.7 57.0 56.6 57.9 49.7 53.2
LLaMA213b 30.0 49.0 34.0 22.5 4.0 38.5 8.5 40.6 10.0 35.4 57.9 46.0 61.9 21.0 30.5 28.0 46.1 23.0 36.1 18.0 26.9 43.0 53.1 55.0 69.4 49.0 33.9 43.1 46.6 42.6
+ CoT 36.0 50.0 38.0 6.0 7.3 39.2 14.0 51.7 10.0 36.9 - 45.0 58.7 30.0 38.9 20.0 40.9 18.0 32.5 21.0 33.6 43.0 58.0 56.0 68.4 47.0 32.5 42.6 47.3 42.4
+ FS 43.0 57.0 60.0 20.5 9.0 46.8 8.0 66.6 15.0 62.3 40.2 24.0 34.2 17.0 18.4 11.0 25.9 5.0 14.6 22.0 33.3 54.0 68.1 50.0 64.8 47.0 45.1 54.0 38.3 43.9
+ FS CoT 37.0 55.0 50.0 33.0 12.0 49.5 11.0 45.6 8.0 44.5 - 35.0 46.0 21.0 25.4 34.0 46.7 23.0 36.5 7.0 16.5 72.0 80.8 54.0 66.2 50.0 43.8 46.5 46.0 45.5
LLaMA27b 39.0 53.0 30.0 13.0 2.7 39.3 4.0 41.0 1.0 6.3 24.5 37.0 49.0 14.0 29.0 7.0 26.8 8.0 21.1 9.0 16.0 48.0 63.9 32.0 47.9 49.0 33.8 27.8 37.8 34.3
+ CoT 44.0 50.0 33.0 5.0 2.7 35.0 4.5 40.0 1.0 1.7 - 27.0 49.9 17.0 31.6 11.0 31.4 10.0 24.5 7.0 17.8 44.0 56.9 32.0 48.1 46.0 33.0 25.6 38.3 34.3
+ FS 44.0 60.0 34.0 11.0 4.0 62.8 8.0 64.7 8.0 40.0 30.5 36.0 50.8 20.0 29.4 5.0 22.3 6.0 18.0 6.0 17.9 12.0 36.3 23.0 44.3 53.0 37.3 49.5 34.0 38.7
+ FS CoT 38.0 51.0 36.0 14.5 11.0 42.8 25.0 65.6 13.0 53.4 - 36.0 53.5 21.0 34.1 1.0 13.6 3.0 11.2 5.0 14.0 22.0 46.7 21.0 42.3 51.0 34.9 53.9 33.3 37.8
Baichuan213b 41.0 61.0 37.0 12.5 4.0 52.0 18.5 63.4 15.0 57.7 52.2 45.0 55.4 29.0 34.6 31.0 48.8 34.0 44.3 30.0 39.5 40.0 57.4 45.0 61.4 49.0 37.9 56.3 48.8 48.0
+ CoT 40.0 57.0 31.0 10.0 3.3 44.6 20.0 61.9 13.0 58.1 - 36.0 41.5 36.0 40.9 39.0 52.0 27.0 38.5 29.0 43.2 46.0 62.8 46.0 64.3 55.0 34.5 54.9 49.8 46.7
+ FS 43.0 59.0 40.0 42.5 24.7 62.1 27.5 70.2 18.0 58.9 63.7 47.0 60.7 35.0 45.7 37.0 51.9 31.0 41.5 19.0 31.8 73.0 81.1 48.0 59.4 48.0 46.1 63.7 52.5 53.7
+ FS CoT 45.0 54.0 48.0 47.0 10.7 44.4 27.0 68.8 15.0 55.0 - 43.0 57.8 27.0 36.7 38.0 49.8 34.0 40.7 33.0 43.0 72.8 80.4 43.0 60.2 44.0 48.5 56.1 51.6 51.7
Baichuan27b 35.0 50.0 37.0 4.5 4.0 47.9 10.5 55.3 15.0 54.3 42.0 26.0 41.5 20.0 34.7 20.0 35.2 19.0 31.2 6.0 20.4 22.0 43.4 29.0 47.7 55.0 31.6 49.9 38.6 39.7
+ CoT 38.0 43.0 32.0 1.0 5.3 37.9 13.0 58.0 15.0 44.2 - 41.0 53.5 28.0 38.8 29.0 39.9 23.0 33.2 18.0 29.3 21.0 41.2 29.0 47.2 54.0 28.5 46.7 42.1 39.4
+ FS 40.0 50.0 36.0 20.0 28.7 59.4 26.5 66.9 17.0 53.0 49.8 45.0 60.7 30.0 42.1 27.0 37.8 23.0 35.7 10.0 20.4 40.0 57.4 37.0 53.0 51.0 36.5 57.3 44.8 45.8
+ FS CoT 41.0 50.0 36.0 23.5 13.0 45.7 17.5 58.1 7.0 39.2 - 36.0 51.2 29.0 43.0 42.0 52.5 25.0 39.3 20.0 31.0 57.0 70.1 39.0 60.2 49.0 37.6 47.7 49.5 46.0
Vicuna1.513b 35.0 50.0 36.0 15.0 8.0 39.2 21.5 59.1 7.0 34.2 51.8 43.0 60.4 29.0 37.0 38.0 46.8 22.0 37.4 17.0 23.2 14.0 42.1 13.0 43.6 46.0 34.0 46.1 42.1 41.1
+ CoT 42.0 51.0 37.0 3.0 1.3 29.8 11.5 50.0 7.0 33.7 - 44.0 56.9 31.0 36.4 16.0 38.2 25.0 37.7 13.0 20.4 31.0 49.0 29.0 49.1 51.0 33.3 37.8 42.3 39.0
+ FS 48.0 57.0 38.0 30.5 7.3 33.6 27.5 57.0 13.0 40.3 39.0 45.0 58.3 23.0 25.9 38.0 42.6 26.0 41.4 18.0 20.1 51.0 61.8 28.0 42.6 56.0 43.4 42.5 43.6 43.3
+ FS CoT 38.0 59.0 39.0 39.5 10.7 37.4 14.0 45.8 12.0 41.6 - 47.0 59.5 27.0 30.7 39.0 48.1 31.0 35.9 26.0 31.2 71.0 77.5 53.0 65.5 52.0 43.9 41.6 50.1 46.7
Vicuna1.57b 37.0 58.0 43.0 5.0 1.3 40.4 9.5 52.5 6.0 32.0 47.8 35.0 47.1 11.0 18.5 20.0 35.7 15.0 25.7 12.0 17.3 14.0 33.0 14.0 46.8 54.0 35.8 43.2 34.8 37.1
+ CoT 36.0 50.0 36.0 1.5 1.3 39.4 8.5 49.2 9.0 36.2 - 30.0 40.9 14.0 24.6 16.0 26.2 14.0 28.5 12.0 25.0 9.0 27.7 7.0 40.3 54.0 30.9 41.6 33.4 34.4
+ FS 43.0 57.0 37.0 8.5 3.3 44.6 5.5 42.1 7.0 36.8 67.1 24.0 31.9 12.0 14.9 16.0 21.8 20.0 27.5 17.0 22.2 13.0 34.3 6.0 32.2 54.0 36.4 47.7 29.9 35.9
+ FS CoT 35.0 54.0 35.0 8.0 2.7 37.2 10.0 47.5 10.0 41.3 - 31.0 39.9 13.0 16.6 15.0 26.7 15.0 23.8 16.0 23.1 55.0 66.3 32.0 48.1 43.0 33.0 42.0 35.9 36.4
FLANT511b 53.0 63.0 43.0 0.0 4.0 52.0 14.0 65.0 13.0 47.7 49.5 56.0 61.7 24.0 26.8 31.0 33.6 48.0 52.2 20.0 21.8 84.0 87.9 78.0 83.9 64.0 39.8 53.6 54.0 50.3
+ CoT 56.0 66.0 45.0 0.0 4.7 49.7 14.5 63.4 13.0 42.7 - 57.6 64.4 23.9 28.2 39.0 41.6 46.0 50.2 28.0 30.6 73.0 79.5 57.0 68.9 55.0 41.8 51.9 52.3 49.4
+ FS 53.0 65.0 43.0 3.5 4.0 50.2 15.5 65.8 9.0 35.0 60.2 55.0 64.3 23.0 25.0 31.0 33.6 47.0 50.6 21.0 22.5 82.0 87.0 78.0 84.5 65.0 41.1 52.8 54.1 50.5
+ FS CoT 54.0 68.0 46.0 3.5 4.0 50.7 13.0 64.0 11.0 43.7 - 54.0 59.4 19.0 21.2 34.0 36.6 45.0 49.2 20.0 21.7 89.0 93.8 72.0 79.7 66.0 42.9 52.8 53.5 50.5
Mistral7b 47.0 50.0 43.0 26.5 11.3 49.8 15.0 58.8 6.0 23.2 58.3 7.0 28.2 5.0 21.4 4.0 24.3 7.0 22.3 4.0 21.7 2.0 39.6 1.0 31.6 51.0 41.6 47.5 30.0 37.3
+ CoT 38.0 56.0 35.0 16.5 13.3 36.6 14.5 49.3 8.0 19.3 - 13.0 31.3 11.0 22.4 8.0 21.1 14.0 24.9 12.0 25.6 5.0 34.0 4.0 31.2 61.0 36.4 35.1 31.4 33.5
+ FS 51.0 57.0 35.0 18.0 12.0 55.8 25.5 71.0 13.0 52.9 63.0 24.0 43.5 12.0 23.9 4.0 21.3 7.0 21.2 7.0 23.0 23.0 48.9 23.0 44.9 57.0 40.3 60.7 35.5 43.0
+ FS CoT 29.0 62.0 32.0 28.5 9.3 46.6 14.5 49.2 8.0 34.4 - 13.0 33.7 11.0 25.3 9.0 26.4 9.0 24.3 9.0 20.3 68.0 78.6 42.0 57.4 50.0 37.9 43.4 39.5 39.8
ChatGLM36b 38.0 50.0 34.0 2.0 3.0 34.1 7.0 43.6 14.0 56.7 38.9 20.0 41.2 14.0 31.7 25.0 33.8 17.0 26.0 24.0 32.2 42.0 57.0 30.0 54.0 50.0 31.0 43.3 40.7 39.0
+ CoT 27.0 49.0 37.0 0.0 1.0 24.8 3.0 37.1 10.0 44.8 - 24.0 41.7 10.0 25.4 27.0 34.6 22.0 28.1 35.0 41.2 28.0 44.5 27.0 52.0 48.0 28.3 35.6 39.4 35.7
+ FS 37.0 52.0 30.0 0.0 10.0 53.0 9.0 52.9 19.0 54.7 50.1 11.0 19.6 13.0 17.0 25.0 35.0 24.0 30.4 20.0 25.6 27.0 46.5 33.0 53.3 54.0 29.8 52.7 35.2 38.2
+ FS CoT 32.0 66.0 31.0 0.0 4.0 34.8 8.0 43.6 11.0 44.0 - 25.0 43.3 19.0 27.5 23.0 30.2 19.0 25.7 36.0 43.0 50.0 62.8 34.0 56.0 48.0 32.3 40.8 42.1 39.2
LLaMA2-Base70b 55.0 61.0 37.0 82.0 40.0 67.4 59.0 85.3 63.0 82.7 74.9 57.0 66.7 36.0 48.3 52.0 61.4 35.0 42.5 25.0 33.8 78.0 85.2 80.0 85.4 61.0 58.8 77.6 60.5 64.4
+ CoT 52.0 73.0 39.0 79.5 32.6 62.3 47.0 79.1 25.0 61.1 55.0 64.3 34.0 43.0 50.0 57.7 37.0 45.2 44.0 53.1 82.0 87.5 76.0 81.6 67.0 60.9 67.5 62.4 63.0
LLaMA2-Base13b 50.0 54.0 30.0 29.5 19.3 53.3 26.5 66.0 20.0 55.6 64.8 48.0 59.3 34.0 48.6 41.0 49.6 38.0 43.4 34.0 37.5 68.0 78.7 49.0 62.7 58.0 40.9 59.9 54.7 52.6
+ CoT 40.0 61.0 37.0 52.0 25.3 59.3 26.0 68.8 11.0 40.8 - 46.0 59.4 37.0 49.1 49.0 58.4 34.0 43.8 38.0 44.1 70.0 78.0 55.0 68.2 58.0 47.5 56.3 57.4 54.5
LLaMA2-Base7b 26.0 50.0 30.0 20.0 19.3 54.5 20.0 59.6 15.0 45.2 62.4 44.0 54.4 30.0 45.3 42.0 49.8 34.0 41.9 30.0 35.8 50.0 64.0 36.0 53.3 49.0 31.5 55.4 49.2 46.3
+ CoT 37.0 52.0 36.0 25.5 21.3 56.9 26.5 67.0 16.0 41.9 - 32.0 45.6 27.0 36.1 41.0 50.9 30.0 38.0 51.0 57.3 45.0 59.7 37.0 57.7 50.0 37.6 55.3 49.4 47.4
Baichuan2-Base13b 38.0 48.0 33.0 42.5 20.7 54.8 42.5 73.0 11.0 45.7 64.9 50.0 59.4 40.0 54.2 42.0 52.7 31.0 38.0 13.0 21.4 68.0 77.3 50.0 63.5 54.0 40.4 59.6 52.6 51.3
+ CoT 50.0 56.0 34.0 47.0 29.3 62.0 22.5 69.3 12.0 43.8 - 46.0 58.2 39.0 49.6 39.0 49.8 34.0 40.1 37.0 45.6 73.0 81.3 46.0 65.6 60.0 46.8 58.4 56.3 54.2
Baichuan2-Base7b 27.0 66.0 41.0 32.5 28.0 59.8 34.5 69.4 5.0 34.3 59.8 40.0 53.8 35.0 50.2 41.0 49.6 33.0 38.5 18.0 22.9 49.0 65.9 34.0 51.0 55.0 41.6 55.8 48.4 48.5
[…]